OpenStudy (anonymous):

A rectangular field is to be subdivided in 7 equal fields. There is 1850 feet of fencing available. Find the dimensions of the field that maximizes the total area. (List the longer side first) Width = __feet Length = __feet What is the maximum area ?

OpenStudy (anonymous):

1850=8W+2L L=925-4W area=L*W area=W*(925-4W) area=925W-4W^2 da/dw (derivative)=925-8W 8W=925 W=115.625 L=1850-4W L=1850-4(115.625) L=1387.5 Max= (1387.5*115.625)=160429.6875

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

Your welcome

OpenStudy (anonymous):

wait..

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

i did all this, and i looked at my work and just had a miscalculation. so then i fixed what i did wrong and i entered it into my hw online but it says it's incorrect.

OpenStudy (anonymous):

what should i do?

OpenStudy (anonymous):

Maybe your homework rounds the length and width to the nearest foot

OpenStudy (anonymous):

oooo ok

OpenStudy (anonymous):

If that does not work let me know

OpenStudy (anonymous):

it's not working

OpenStudy (anonymous):

Does it say all three are incorrect or just one?

OpenStudy (anonymous):

the length says that it is 115.625. everything else is incorrect

OpenStudy (anonymous):

Oh I know why L=925-4W L=462.5 Area= 53476.6525

OpenStudy (anonymous):

Before I had calculated L=1850-4W.. my apologies

OpenStudy (anonymous):

no no worries! thank you very much! =)

OpenStudy (anonymous):

No problem, have a nice day. Bye.